# Complexity of codes for Ramsey positive sets

### v1

@@ 1. Introduction
Sabok proved that the set of codes for $G_{\delta}$ (i.e., $\mathbf{\Pi}^{0}_{2}$) Ramsey positive sets in $[\omega]^{\omega}$ is $\mathbf{\Sigma}^{1}_{2}$-complete [Sab12, Theorem 1]; see Section 3 for the relevant definitions. In the same paper, Sabok proved a general theorem about $\mathbf{\Sigma}^{1}_{2}$-complete sets that has been used in several recent proofs of $\mathbf{\Sigma}^{1}_{2}$-completeness [Sab12, Theorem 2]. For instance, Todorčević and Vidnyánszky showed that the set of closed subgraphs of the shift graph on $[\omega]^{\omega}$ that have finite Borel chromatic number is $\mathbf{\Sigma}^{1}_{2}$-complete [TV21, Theorem 1.3]. Brandt, Chang, Grebík, Grunau, Rozhoň, and Vidnyánszky showed a similar result for graphs of bounded degree: for any $d>2$, the set of Borel acyclic $d$-regular graphs with Borel chromatic number at most $d$ is $\mathbf{\Sigma}^{1}_{2}$-complete [BCG+21, Theorem 1.2]. Finally, Thornton proved that a more general class of Borel constraint satisfaction problems is $\mathbf{\Sigma}^{1}_{2}$-complete, including several other examples from Borel combinatorics [Tho22, Theorem 1.7].
To prove [Sab12, Theorem 1], Sabok considered the set of codes corresponding to a universal $G_{\delta}$ set formed by viewing every $x\in 2^{\omega}$ as representing a countable sequence of trees. Under certain conditions, we show that we can find a continuous reduction from this set of codes to the analogous set of codes associated with a topological Ramsey space satisfying axioms specified by Todorčević in [Tod10]. In particular,
@@
Suppose $(\mathcal{R},\leq,r)$ satisfies A.1-A.4 from [Tod10, Section 5.1], $\mathcal{R}$ is closed, $\mathcal{AR}$ is countable, and there is a function $m:\mathcal{AR}\to\omega$ satisfying Property 5.1 (see Section 5). Then the set of codes for $G_{\delta}$ Ramsey positive subsets of $\mathcal{R}$ is $\mathbf{\Sigma}^{1}_{2}$-complete.
@@ 2. Acknowledgments
Thank you to Clinton Conley, Natasha Dobrinen, and Aristotelis Panagiotopoulos for helpful discussions. The author was supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE2140739, and by the ARCS Foundation.
@@ 3. Background
Throughout, we will let $(\mathcal{R},\leq,r)$ denote a triple where $\mathcal{R}$ is a non-empty set, $\leq$ is a reflexive and transitive relation on $\mathcal{R}$, and $r:\mathcal{R}\times\omega\to\mathcal{AR}$ is a map into a set of finite approximations of $\mathcal{R}$. We assume that $\mathcal{AR}$ is countable. In addition, we assume that $(\mathcal{R},\leq,r)$ satisfies axioms A.1, A.2, A.3, and A.4 from [Tod10, Section 5.1] and that $\mathcal{R}$ is closed as a subspace of $\mathcal{AR}^{\omega}$, where $\mathcal{AR}$ is given the discrete topology. Unless otherwise stated, we use $s,t,\ldots$ to denote elements of $\mathcal{AR}$ and $A,B,\ldots$ to denote elements of $\mathcal{R}$.
@@
Under the assumptions above, every subset of $\mathcal{R}$ with the Baire property in the Ellentuck topology is Ramsey, and every Ellentuck meager set is Ramsey null.
Since $\mathcal{R}$ is Polish, every analytic subset of $\mathcal{R}$ is Suslin-measurable, hence Ramsey by [Tod10, Corollary 5.9]. Then for every analytic $\mathcal{X}\subseteq\mathcal{R}$, we have that $\mathcal{X}$ is Ramsey positive iff $\mathcal{X}$ contains some $[s,A]\neq\emptyset$.
@@ 4. The set of codes for $G_{\delta}$ Ramsey positive sets is $\mathbf{\Sigma}^{1}_{2}$
Note that condition (1) is closed, so it suffices to verify that condition (2) is $\mathbf{\Sigma}^{1}_{2}$. Since $\mathcal{R}\setminus\bigcup_{n<\omega}[X_{n}]$ is analytic, condition (2) holds iff $\mathcal{R}\setminus\bigcup_{n<\omega}[X_{n}]$ contains some non-empty $[s,A]$, i.e.,
@@ 5. Conditions for $\mathbf{\Sigma}^{1}_{2}$-completeness
Property 5.1.
We consider the following four properties that a map $m:\mathcal{AR}\to\omega$ may satisfy:
@@
    (Compatibility with $\leq$) There exists $A^{*}$ such that ¯$\overline{\ell}(A^{*})=\omega$ and for all $B\leq A\leq A^{*}$, we have ¯$\overline{\ell}(B)\subseteq\overline{\ell}(A)$.
@@
Remark 5.2.
Instead of requiring ¯$\overline{\ell}(A^{*})=\omega$ in the compatibility property, it is sufficient for ¯$\overline{\ell}(A^{*})$ to be maximal in the sense that for every $A\in\mathcal{R}$, we have ¯$\overline{\ell}(A)\subseteq\overline{\ell}(A^{*})$. Given a map $m$ satisfying this modified version of compatibility along with monotonicity, unboundedness, and selection, a slight adjustment will yield a map satisfying all of Property 5.1.
@@
Note that monotonicity implies $\ell(s)\sqsubseteq\ell(t)$ when $s\sqsubseteq t$, and $\ell(s)\sqsubseteq\overline{\ell}(A)$ when $s\sqsubseteq A$. In addition, Property 5.1 implies that $\ell[\mathcal{AR}]=[\omega]^{<\omega}$ and ¯$\overline{\ell}[\mathcal{R}]=[\omega]^{\omega}$; compatibility and selection are key to this observation.
@@
Suppose there exists a map $m:\mathcal{AR}\to\omega$ satisfying Property 5.1. Then $C$ is $\mathbf{\Sigma}^{1}_{2}$-hard.
@@ 6. Examples
We present a few examples of topological Ramsey spaces where Theorem 1.1 shows that the set of codes for $G_{\delta}$ Ramsey positive sets is $\mathbf{\Sigma}^{1}_{2}$-complete. For each of these spaces, we exhibit a map $m$ witnessing Property 5.1.
@@ 6.1. Ellentuck space
The proof of Proposition 5.4 relies on Sabok’s result that the set of codes for $G_{\delta}$ Ramsey positive sets of $[\omega]^{\omega}$ is $\mathbf{\Sigma}^{1}_{2}$-complete, so finding a map $m:[\omega]^{<\omega}\to\omega$ satisfying Property 5.1 is superfluous. We nevertheless provide such a map for the sake of illustration.
Define $m:[\omega]^{<\omega}\to\omega$ by $m(\emptyset):=0$ and $m(s):=\max(s)$ for $s\neq\emptyset$. Then $\ell(s)=s$ and ¯$\overline{\ell}(A)=A$ for all $s\in[\omega]^{<\omega}$ and $A\in[\omega]^{\omega}$. It is clear that monotonicity and unboundedness hold. Letting $A^{*}=\omega$, we see that compatibility holds as well. Finally, selection holds since we can pick $B=B^{\prime}$. So $m$ satisfies Property 5.1.
@@ 6.2. Strong subtrees of $2^{<\omega}$
Thus, $m$ satisfies Property 5.1, and Theorem 1.1 implies that the set of codes for $G_{\delta}$ Ramsey positive subsets of $\mathcal{S}_{\infty}(2^{<\omega})$ is $\mathbf{\Sigma}^{1}_{2}$-complete.
@@ 6.3. Infinite sequences of words with variables
We conclude that $m$ satisfies Property 5.1, hence the set of codes for $G_{\delta}$ Ramsey positive subsets of $W_{Lv}^{[<\infty]}$ is $\mathbf{\Sigma}^{1}_{2}$-complete by Theorem 1.1.
@@ 6.4. High-dimensional Ellentuck spaces
For $k\geq 2$, we define the topological Ramsey space $\mathcal{E}_{k}$ as in [Dob16a]. Let $\omega^{\not\downarrow\leq k}$ denote the set of all non-decreasing sequences of natural numbers of length at most $k$, i.e.,
@@
Note that condition (ii) guarantees that every $A\in\mathcal{E}_{k}$ is uniquely determined by $A[\omega^{\not\downarrow k}]=\{A(\vec{i}_{n})\mid n<\omega\}$. Furthermore, there is a unique $\mathcal{E}_{k}$-tree $\widehat{A}$ for which $[\widehat{A}]=A$ by conditions (i) and (iii). For $A,B\in\mathcal{E}_{k}$, we define $A\leq B$ iff $A[\omega^{\not\downarrow k}]\subseteq B[\omega^{\not\downarrow k}].$ For $n<\omega$ and $A\in\mathcal{E}_{k}$, let $r_{n}(A):=A\mathord{\upharpoonright}\{\vec{i}_{p}\mid p<n\}$. By [Dob16a, Theorem 3.17], $(\mathcal{E}_{k},\leq,r)$ is a topological Ramsey space.
@@
We check that $m$ satisfies Property 5.1. To see monotonicity, note that if $\emptyset\neq s\sqsubseteq t$, then we have $M_{s}\leq M_{t}$, $\langle M_{s}\rangle\preceq\langle M_{t}\rangle$, and $s(\vec{i}_{n_{s}})=t(\vec{i}_{n_{s}})$. Fix any $A\in\mathcal{E}_{k}$ with $t\sqsubseteq A$. Then
@@
It is not difficult to see that ¯$\overline{\ell}(A)=\{\max(\widehat{A}(n))\mid n<\omega\}$ for all $A\in\mathcal{E}_{k}$. Since $\max(\widehat{A}(0))<\max(\widehat{A}(1))<\ldots$ by condition (ii), we have ¯$\overline{\ell}(A)\in[\omega]^{\omega}$, hence unboundedness is satisfied.
Observe that $m$ does not satisfy compatibility as written. We will instead show that $m$ satisfies the modified version of compatibility given in Remark 5.2. If we set $A^{*}=W_{k}$, then $m(A^{*})$ is maximal. It suffices to verify that $B\leq A\leq A^{*}$ implies ¯$\overline{\ell}(B)\subseteq\overline{\ell}(A)$. Note that
@@
whenever $B\leq A$, so $m$ satisfies the modified version of compatibility.
@@
Thus, $m$ satisfies Property 5.1 with the modification of compatibility according to Remark 5.2, hence Theorem 1.1 applies.
@@ 6.5. Other examples
We note that we can find a map satisfying Property 5.1 for the following spaces: $\text{FIN}_{k}^{[\infty]}$ as defined in [Tod10, Section 5.2]; $\mathcal{E}_{\infty}$ as defined in [Tod10, Section 5.6]; $\text{FIN}_{Lv}^{[\infty]}$ as defined in [Tod10, Section 5.3]; $\mathcal{M}_{\infty}$ as defined in [Tod10, Section 5.7]; $\mathcal{R}_{1}$ as defined in [DT14]. So Theorem 1.1 applies to all of these spaces.
There are several generalizations of spaces mentioned above where we have not been able to verify the existence of a map satisfying Property 5.1. In [Dob16b], Dobrinen defined spaces $\mathcal{E}_{B}$ for uniform barriers on $\omega$ as an extension of the spaces $\mathcal{E}_{k}$ from [Dob16a]. We suspect that a map similar to the $m$ defined in Section 6.4 may satisfy Property 5.1 in $\mathcal{E}_{B}$; however, we have not been able to prove it. Likewise, we do not know if we can apply Theorem 1.1 to the spaces $\mathcal{R}_{\alpha}$, $\alpha<\omega_{1}$, defined in [DT15] as an extension of the space $\mathcal{R}_{1}$ defined in [DT14].

### v2

@@ 1. Introduction
Sabok proved that the set of codes for $G_{\delta}$ (i.e., $\mathbf{\Pi}^{0}_{2}$) Ramsey positive sets in $[\omega]^{\omega}$ is $\mathbf{\Sigma}^{1}_{2}$-complete [Sab12, Theorem 1]; see Section 3 for the relevant definitions. In the same paper, Sabok proved a general theorem about $\mathbf{\Sigma}^{1}_{2}$-complete sets that has been used in several recent proofs of $\mathbf{\Sigma}^{1}_{2}$-completeness [Sab12, Theorem 2]. For instance, Todorčević and Vidnyánszky showed that the set of closed subgraphs of the shift graph on $[\omega]^{\omega}$ that have finite Borel chromatic number is $\mathbf{\Sigma}^{1}_{2}$-complete [TV21, Theorem 1.3]. Brandt, Chang, Grebík, Grunau, Rozhoň, and Vidnyánszky showed a similar result for graphs of bounded degree: for any $d>2$, the set of Borel acyclic $d$-regular graphs with Borel chromatic number at most $d$ is $\mathbf{\Sigma}^{1}_{2}$-complete [BCG+24, Theorem 1.2]. Finally, Thornton proved that a more general class of Borel constraint satisfaction problems is $\mathbf{\Sigma}^{1}_{2}$-complete, including several other examples from Borel combinatorics [Tho22, Theorem 1.7].
To prove [Sab12, Theorem 1], Sabok considered the set of codes corresponding to a universal $G_{\delta}$ set formed by viewing every $x\in 2^{\omega}$ as representing a countable sequence of trees. We show that we can find a continuous reduction from this set of codes to the analogous set of codes associated with a topological Ramsey space satisfying axioms specified by Todorčević in [Tod10] whenever the space is sufficiently similar to $[\omega]^{\omega}$. We call such spaces well-indexed; see Definition 5.1.
@@
Suppose $(\mathcal{R},\leq,r)$ satisfies A.1-A.4 from [Tod10, Section 5.1], $\mathcal{R}$ is closed, $\mathcal{AR}$ is countable, and $(\mathcal{R},\leq,r)$ is well-indexed. Then the set of codes for $G_{\delta}$ Ramsey positive subsets of $\mathcal{R}$ is $\mathbf{\Sigma}^{1}_{2}$-complete.
@@ 2. Acknowledgments
Thank you to Clinton Conley, Natasha Dobrinen, and Aristotelis Panagiotopoulos for helpful discussions, and to the referee for many useful suggestions. The author was supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE2140739, and by the ARCS Foundation.
@@ 3. Background
Throughout, we will let $(\mathcal{R},\leq,r)$ denote a triple where $\mathcal{R}$ is a non-empty set, $\leq$ is a reflexive and transitive relation on $\mathcal{R}$, and $r:\mathcal{R}\times\omega\to\mathcal{AR}$ is a map into a set of finite approximations of $\mathcal{R}$. We assume that $\mathcal{AR}$ is countable. In addition, we assume that $(\mathcal{R},\leq,r)$ satisfies axioms A.1, A.2, A.3, and A.4 from [Tod10, Section 5.1] and that $\mathcal{R}$ is closed as a subspace of $\mathcal{AR}^{\omega}$, where $\mathcal{AR}$ is given the discrete topology. Unless otherwise stated, we use $s,t,\ldots$ to denote elements of $\mathcal{AR}$ and $A,B,\ldots$ to denote elements of $\mathcal{R}$. We write $\mathcal{R}$ for the triple $(\mathcal{R},\leq,r)$ when clear from context.
@@
If $(\mathcal{R},\leq,r)$ satisfies axioms A.1-A.4 from [Tod10, Section 5.1] and $\mathcal{R}$ is closed as a subset of $\mathcal{AR}^{\omega}$, then every subset of $\mathcal{R}$ with the Baire property in the Ellentuck topology is Ramsey, and every Ellentuck meager set is Ramsey null.
Every analytic set has the Baire property in the Ellentuck topology and is therefore Ramsey [Ell74, Corollary 11]. So for every analytic $\mathcal{X}\subseteq\mathcal{R}$, we have that $\mathcal{X}$ is Ramsey positive iff $\mathcal{X}$ contains some $[s,A]\neq\emptyset$.
@@ 4. The set of codes for $G_{\delta}$ Ramsey positive sets is $\mathbf{\Sigma}^{1}_{2}$
Note that condition (1) is closed, so it suffices to verify that condition (2) is $\mathbf{\Sigma}^{1}_{2}$. Since $\mathcal{R}\setminus\bigcup_{n<\omega}[X_{n}]$ is analytic, condition (2) holds iff $\mathcal{R}\setminus\bigcup_{n<\omega}[X_{n}]$ contains some non-empty $[s,A]$; equivalently,
@@ 5. Conditions for $\mathbf{\Sigma}^{1}_{2}$-completeness
Definition 5.1.
A topological Ramsey space $(\mathcal{R},\leq,r)$ is well-indexed if there is a map $m:\mathcal{AR}\to\omega$ satisfying the following properties:
@@
    (Compatibility with $\leq$) There exists $A^{*}$ such that ¯$\overline{\ell}(A^{*})$ is maximal (i.e., ¯$\overline{\ell}(A)\subseteq\overline{\ell}(A^{*})$ for all $A$) and for all $B\leq A\leq A^{*}$, we have ¯$\overline{\ell}(B)\subseteq\overline{\ell}(A)$.
@@
Suppose $m:\mathcal{AR}\to\omega$ well-indexes $\mathcal{R}$ and $A^{*}$ witnesses that $m$ satisfies compatibility. Note that monotonicity implies $\ell(s)\sqsubseteq\ell(t)$ when $s\sqsubseteq t$, and $\ell(s)\sqsubseteq\overline{\ell}(A)$ when $s\sqsubseteq A$. In addition, we have $\ell[\mathcal{AR}]=[\overline{\ell}(A^{*})]^{<\omega}$ and ¯[]=[¯$\overline{\ell}[\mathcal{R}]=[\overline{\ell}(A^{*})]^{\omega}$; compatibility and selection are key to this observation.
@@
Suppose $\mathcal{R}$ is well-indexed. Then $C$ is $\mathbf{\Sigma}^{1}_{2}$-hard.
@@
Let $m:\mathcal{AR}\to\omega$ well-index $\mathcal{R}$. We may assume that for $A^{*}$ witnessing compatibility, ¯$\overline{\ell}(A^{*})=\omega$: if not, let $g:\overline{\ell}(A^{*})\to\omega$ be an order-preserving bijection, and consider $m^{\prime}:\mathcal{AR}\to\omega$ defined by $m^{\prime}(\emptyset):=0$ and $m^{\prime}(s):=g(m(s))$ if $s\neq\emptyset$. Then $m^{\prime}$ well-indexes $\mathcal{R}$, and our original $A^{*}$ witnesses compatibility with ¯$\overline{\ell^{\prime}}(A^{*})=\omega$.
@@ 6. Examples
We present a few examples of topological Ramsey spaces where Theorem 1.1 shows that the set of codes for $G_{\delta}$ Ramsey positive sets is $\mathbf{\Sigma}^{1}_{2}$-complete. For each, we exhibit a map $m$ that well-indexes the space.
@@ 6.1. Ellentuck space
The proof of Proposition 5.3 relies on Sabok’s result that the set of codes for $G_{\delta}$ Ramsey positive sets of $[\omega]^{\omega}$ is $\mathbf{\Sigma}^{1}_{2}$-complete, so showing that $[\omega]^{\omega}$ is well-indexed is superfluous. We nevertheless provide such a map for the sake of illustration.
Define $m:[\omega]^{<\omega}\to\omega$ by $m(\emptyset):=0$ and $m(s):=\max(s)$ for $s\neq\emptyset$. Then $\ell(s)=s$ and ¯$\overline{\ell}(A)=A$ for all $s\in[\omega]^{<\omega}$ and $A\in[\omega]^{\omega}$. It is clear that monotonicity and unboundedness hold. Letting $A^{*}=\omega$, we see that compatibility holds as well. Finally, selection holds since we can pick $B=B^{\prime}$. So $m$ well-indexes $[\omega]^{\omega}$.
@@ 6.2. Strong subtrees of $2^{<\omega}$
Thus, $\mathcal{S}_{\infty}(2^{<\omega})$ is well-indexed, and Theorem 1.1 implies that the set of codes for $G_{\delta}$ Ramsey positive subsets of $\mathcal{S}_{\infty}(2^{<\omega})$ is $\mathbf{\Sigma}^{1}_{2}$-complete.
@@ 6.3. Infinite sequences of words with variables
We conclude that $W_{Lv}^{[<\infty]}$ is well-indexed, hence the set of codes for $G_{\delta}$ Ramsey positive subsets of $W_{Lv}^{[<\infty]}$ is $\mathbf{\Sigma}^{1}_{2}$-complete by Theorem 1.1.
@@ 6.4. High-dimensional Ellentuck spaces
For $k\geq 2$, we define the topological Ramsey space $\mathcal{E}_{k}$ as in [Dob16a]. Let $\omega^{\not\downarrow\leq k}$ denote the set of all non-decreasing sequences of natural numbers of length at most $k$, that is,
@@
Note that condition (2) guarantees that every $A\in\mathcal{E}_{k}$ is uniquely determined by $A[\omega^{\not\downarrow k}]=\{A(\vec{i}_{n})\mid n<\omega\}$. Furthermore, there is a unique $\mathcal{E}_{k}$-tree $\widehat{A}$ for which $[\widehat{A}]=A$ by conditions (1) and (3). For $A,B\in\mathcal{E}_{k}$, we define $A\leq B$ iff $A[\omega^{\not\downarrow k}]\subseteq B[\omega^{\not\downarrow k}].$ For $n<\omega$ and $A\in\mathcal{E}_{k}$, let $r_{n}(A):=A\mathord{\upharpoonright}\{\vec{i}_{p}\mid p<n\}$. By [Dob16a, Theorem 3.17], $(\mathcal{E}_{k},\leq,r)$ is a topological Ramsey space.
@@
We check that $m$ well-indexes $\mathcal{E}_{k}$. To see monotonicity, note that if $\emptyset\neq s\sqsubseteq t$, then we have $M_{s}\leq M_{t}$, $\langle M_{s}\rangle\preceq\langle M_{t}\rangle$, and $s(\vec{i}_{n_{s}})=t(\vec{i}_{n_{s}})$. Fix any $A\in\mathcal{E}_{k}$ with $t\sqsubseteq A$. Then
@@
It is not difficult to see that ¯$\overline{\ell}(A)=\{\max(\widehat{A}(n))\mid n<\omega\}$ for all $A\in\mathcal{E}_{k}$. Since $\max(\widehat{A}(0))<\max(\widehat{A}(1))<\ldots$ by condition (2), we have ¯$\overline{\ell}(A)\in[\omega]^{\omega}$, hence unboundedness is satisfied.
We next verify compatibility. If we set $A^{*}=W_{k}$, then $m(A^{*})$ is maximal. It suffices to verify that $B\leq A\leq A^{*}$ implies ¯$\overline{\ell}(B)\subseteq\overline{\ell}(A)$. Note that
@@
whenever $B\leq A$, so $m$ satisfies compatibility.
@@
Thus, $\mathcal{E}_{k}$ is well-indexed, hence Theorem 1.1 applies.
@@ 6.5. Other examples
We provide several other examples to which Theorem 1.1 applies. We exhibit a map $m$ well-indexing each space, using the notation from the indicated reference. The verification that these functions satisfy the required properties is omitted.
  1. (1)
    $\text{FIN}_{k}^{[\infty]}$ as defined in [Tod10, Section 5.2]: Define $m:\text{FIN}_{k}^{[<\infty]}\to\omega$ by $m(\emptyset):=0$ and $m((p_{i})_{i<n}):=\min\{j\mid p_{n-1}(j)=k\}$ for $n\geq 1$.
  2. (2)
    $\text{FIN}_{Lv}^{[\infty]}$ as defined in [Tod10, Section 5.3]: Define $m:\text{FIN}_{Lv}^{[<\infty]}\to\omega$ by $m(\emptyset):=0$ and $m((x_{i})_{i<n}):=\min\{j\mid x_{n-1}(j)=v\}$ for $n\geq 1$.
  3. (3)
    $\mathcal{E}_{\infty}$ as defined in [Tod10, Section 5.6]: Define $m:\mathcal{AE}_{\infty}\to\omega$ by $m(\emptyset):=0$ and $m(r_{n}(E)):=p_{n}(E)$ if $n\geq 1$.
  4. (4)
    $\mathcal{M}_{\infty}$ as defined in [Tod10, Section 5.7]: Define $m:\mathcal{AM}_{\infty}\to\omega$ by $m(\emptyset):=0$ and $m(r_{n}(A)):=p_{n-1}(A)$ if $n\geq 1$.
  5. (5)
    $\mathcal{R}_{1}$ as defined in [DT14]: Define $m:\mathcal{AR}\to\omega$ by $m(\emptyset):=0$ and $m(a):=k_{n-1}$ for $a\in\mathcal{AR}_{n}$ with $n\geq 1$ and $a(i)\subseteq\mathbb{T}(k_{i})$ for each $i<n$.
There are several generalizations of spaces mentioned above that we have not been able to prove are well-indexed. In [Dob16b], Dobrinen defined spaces $\mathcal{E}_{B}$ for uniform barriers on $\omega$ as an extension of the spaces $\mathcal{E}_{k}$ from [Dob16a]. We suspect that a map similar to the $m$ defined in Section 6.4 may well-index $\mathcal{E}_{B}$; however, we have not been able to prove it. Likewise, we do not know if we can apply Theorem 1.1 to the spaces $\mathcal{R}_{\alpha}$, $\alpha<\omega_{1}$, defined in [DT15] as an extension of the space $\mathcal{R}_{1}$ defined in [DT14].